If R1=11110000, R2=00001111 what is the value of R1 after the execution of R1 - R1+R2? Answer: on page
Answer #1

R1=11110000

R2=00001111

After the exectution of R1<-- R1+R2, the value of R1 is find below:

As we know, the meaning of this execution (R1<-- R1+R2) is addition the value of R1 and R2 and then transfer value in R1.

Binary Addition rules:

1 + 0 = 1

0 + 1 = 1

0 + 0 = 0

1 + 1 = 10

So, the addition of R1 and R2 is

R1 = 11110000 R2 = 00001111 + 11111111So, after the exection of R1<-- R1+R2. The value of R1 is: 11111111